### v1

# Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3

Abstract
As the parameters of large language models (LLMs) expand, the computational cost of fine-tuning the entire model becomes prohibitive. To address this challenge, we introduce a parameter-efficient fine-tuning (PEFT) method, Principal Singular values and Singular vectors Adaptation (PiSSA), which optimizes a significantly reduced parameter space while achieving or surpassing the performance of full-parameter fine-tuning. PiSSA is inspired by Intrinsic SAID, which suggests that pre-trained, over-parametrized models inhabit a space of low intrinsic dimension. Consequently, PiSSA represents a matrix $W\in\mathbb{R}^{m\times n}$ within the model by the product of two trainable matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, plus a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ for error correction. Singular value decomposition (SVD) is employed to factorize $W$, and the principal singular values and vectors of $W$ are utilized to initialize $A$ and $B$. The residual singular values and vectors initialize the residual matrix $W^{res}$, which keeps frozen during fine-tuning. Notably, PiSSA shares the same architecture with Low-Rank Adaptation (LoRA), which hypothesizes that changes in model parameters $\Delta W$ form a low-rank matrix. However, LoRA approximates $\Delta W$ through the product of two matrices, $A$, initialized with Gaussian noise, and $B$, initialized with zeros, while PiSSA initializes $A$ and $B$ with principal singular values and singular vectors of the original matrix $W$. Given that the principal singular values and vectors capture the essence of a low-rank matrix, PiSSA can better approximate the outcomes of full-parameter fine-tuning at the beginning by changing the essential parts while freezing the “noisy” parts. In comparison, LoRA freezes the original matrix and updates the “noise”. This distinction enables PiSSA to convergence much faster than LoRA and also achieve better performance in the end. On five common benchmarks, PiSSA outperforms LoRA on all of them using exactly the same setups except for a different initialization. On GSM8K, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, outperforming LoRA’s 67.7% by 5.16%. Due to the same architecture, PiSSA inherits many of LoRA’s advantages, such as parameter efficiency and compatibility with quantization. Leveraging a fast SVD method, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
1 Introduction
Fine-tuning large language models (LLMs) is a highly effective technique for boosting their capabilities in various tasks [1, 2, 3, 4], ensuring models to follow instructions [5, 6, 7], and instilling models with desirable behaviors while eliminating undesirable ones [8, 9]. However, the fine-tuning process for very large models is accompanied by prohibitive costs. For example, regular 16-bit fine-tuning of a LLaMA 65B parameter model requires over 780 GB of GPU memory [10], and the VRAM consumption for training GPT-3 175B reaches 1.2TB [11]. Consequently, various parameter-efficient fine-tuning (PEFT) [12, 13] methods have been proposed to reduce the number of parameters and memory usage required for fine-tuning. Due to the ability to maintain the performance of full fine-tuning without adding additional inference latency, Low-Rank Adaptation (LoRA) [11] has emerged as a popular PEFT method.
LoRA [11] hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties. As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$. For $h=Wx$, the modified forward pass computation is expressed as follows:
$h=(W+\Delta W)x=(W+AB)x.$ (1)
A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix, instead optimizing the injected, significantly smaller low-rank matrices. Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times. Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks. By integrating the quantization of pre-trained models, LoRA also enables reducing the average memory requirements by 16 times [10]. Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full precision base models.
(a) Full Fine-tuning
(b) LoRA
(c) PiSSA
Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA. In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates.
Inspired by the core idea of Intrinsic SAID [14, 15], our Principal Singular values and Singular vectors Adapter (PiSSA) diverges from LoRA and its successors by not focusing on approximating $\Delta W$, but $W$. In alignment with Intrinsic SAID, we assume that the matrix $W$ in pre-trained models exhibits low (intrinsic) rank characteristics. Thus, we apply singular value decomposition (SVD) to matrix $W$. Based on the magnitude of the singular values, we divide $W$ into two components: the principal matrix $W^{pri}$ and the residual matrix $W^{res}$. The principal matrix $W^{pri}$ is of low rank and can be represented by the product of $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$. $A$ and $B$ are initialized based on the principal singular values and vectors and are trainable, whereas $W^{res}$ is initialized with the product of the residual singular values and vectors and are frozen. As indicated by Equation (2), the integration of $AB$ (the PiSSA adaptor) with the residual model also preserves the capability of the pre-trained model in the beginning of fine-tuning:
$h=Wx=(W^{res}+W^{pri})x=(W^{res}+AB)x.$ (2)
Different from LoRA, a significantly less important part $W^{res}$ is kept frozen, while the essential part $W^{pri}$ is fully tunable from the beginning, which enables PiSSA to fit the training data faster and better.
2 Related Works
The vast complexity and computational needs of large language models (LLMs) with billions of parameters present significant hurdles in adapting them for specific downstream tasks. Parameter Efficient Fine-Tuning (PEFT) [12, 13] emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning. PEFT encompasses strategies like partial fine-tuning, soft prompt fine-tuning, non-linear adapter fine-tuning, and linear adapter fine-tuning.
Partial fine-tuning selects certain model parts, e.g., bias and layer normalization [16, 17], or inserted trainable mask to the pre-trained weights [18, 19] and delta weights [20, 21, 22, 23], for fine-tuning. However, the performance of partial fine-tuning may be limited due to only specific layers’ parameters being fine-tuned. Soft prompt fine-tuning  [24, 25, 26, 27, 28, 29, 30] introduces optimizable vectors into the model’s input or hidden layers, adapting to specific tasks with minimal structural changes. Non-linear adapter fine-tuning  [31, 32, 33, 34, 35, 36, 37, 38, 39, 40] incorporates (in a non-linear manner) trainable (non-linear) modules to the models. Both soft prompt fine-tuning and non-linear adapter fine-tuning add complexity and inference costs despite their adaptability. Linear adapter fine-tuning [14, 15, 11, 41], particularly LoRA [11] and its derivative methods, involves adding trainable parameters to the self-attention and feed-forward layers, which are then linearly combined with the original weights. After fine-tuning, these adaptations can be reparametrized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning.
LoRA [11] posits that weight adjustments during model adaptation exhibit a low "intrinsic rank". It learns the weight updates through a low-rank decomposition, keeping the original structure frozen and proving effective in parameter-efficient fine-tuning. Following LoRA, Low-rank Adjustment [42, 41, 43] learns the rank size needed for LoRA in each layer of the model. LoRA-guided Pretrained Weight Update [44, 45] updates the original weights of the model using the parameters from adapter layers, enhancing LoRA’s representational capacity. Quantization Adaptation [10, 46, 47] applies quantization to the base model, further reducing training consumption overhead.
Differing from LoRA’s approach of learning low-rank approximations of weight updates, our PiSSA method innovates by utilizing the low intrinsic dimension of pre-trained weight directly. It initializes updatable adapter layers with principal singular values and vectors from the model’s weights, preserving the essential capabilities while enabling effective fine-tuning. This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like Low-rank Adjustment for enhanced fine-tuning efficiency.
3 Method
[…]
$A$ $=U_{[:r]}\,\text{diag}(S_{[:r]}^{1/2})\in\mathbb{R}^{m\times r},$ (3)
$B$ $=\text{diag}(S_{[:r]}^{1/2})\,V_{[:r]}^{T}\in\mathbb{R}^{r\times n}.$ (4)
The residual singular values and vectors are used to build the residual matrix:
$W^{res}=U_{[r:]}\,\text{diag}(S_{[r:]})\,V_{[r:]}^{T}\in\mathbb{R}^{m\times n}.$ (5)
Given $W=W^{res}+AB$, the introduction of PiSSA in the beginning of training does not compromise the model’s capabilities. Considering $S_{[r:]}\approx 0$, thus rendering $W^{res}\approx 0$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities. Then, the fine-tuning of $AB$ mirrors the process of fine-tuning the entire model, despite using much less parameters. In contrast, LoRA initializes the adapter with Gaussian noise and zeros, thus the principal part of $W$ is not updatable in the early stage of the fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include but not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, the quantization of the residual model to decrease memory consumption during forward propagation in training, and easy deployment. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\{A,B\}$ instead. When we want to utilize a fine-tuned PiSSA, similar to initializing one, we first perform singular value decomposition on the base model (which is very fast as will be shown in our experiments) and then assign $\{A,B\}$ to the adapter. Finally, one pre-trained model can accommodate multiple $\{A,B\}$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.
4 Experiments
We conduct comparative experiments between PiSSA, LoRA, and full parameter fine-tuning on LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50]. These comparing cover a wide range of tasks, including math [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6]. The findings reveal that fine-tuning adapters initialized with principal singular values and vectors (PiSSA) outperform those initialized with intermediate or less significant ones. This evidences that direct fine-tuning of the model’s principal components leads to superior outcomes (Section 4.1). PiSSA demonstrates superior performance on downstream tasks (Section 4.3). Subsequently, we conduct experiments across various ranks, observing that PiSSA converges more swiftly and aligns more closely with the training data compared to LoRA, exhibiting robust superiority under the same trainable parameter configurations (Section 4.2). Finally, we employ subsequent experiments to demonstrate that the Fast SVD [56] technique helps our PiSSA achieve a balance between initialization speed and performance (Section 4.4).
In our experiments, we adopt the Alpaca [57] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. We utilize the bf16 computation type, ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model. We train only on the first 100,000 conversions from the dataset for one epoch, and only using the response of the conversions to compute loss. The experiments in this section were conducted on the NVIDIA A800-SXM4(80G) GPU.
4.1 Verifying the importance of principal singular values and vectors
To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors. These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [51] and MATH datasets [52], with the outcomes depicted in Figures 2.
Figure 2: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models. This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
4.2 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
This section explores the impact of incrementally increasing the rank of PiSSA and LoRA from 1 to 128, aiming to determine the optimal number of trainable parameters needed to achieve performance comparable to full parameter training. The training is conducted using the MetaMathQA dataset, while the validation is performed on the GSM8K and MATH datasets. The outcomes of these experiments are depicted in Figure 3, with additional results presented in Figure 4 in the Appendix.
(a) Variation of loss with respect to rank 1 throughout the training phase. Additional ranks are depicted in Figure 4 in the Appendix.
(b) Final training loss across different ranks.
(c) Rank-wise performance evaluated using pass@1 on the GSM8K dataset.
(d) Rank-wise performance evaluated using pass@1 on the MATH dataset.
Figure 3: Fine-tuning LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
From Figure 3(a), it is observed that at the beginning of training, PiSSA quickly starts to converge, whereas LoRA experiences a period where the loss not only fails to decrease but even slightly increases. Furthermore, throughout the entire training process, the loss of PiSSA consistently remains at a lower level compared to LoRA. Figure 3(b) shows the final loss on the training set for models trained with ranks varying from 1 to 128. The data indicate that, compared to LoRA, PiSSA achieves a level of fitting to the training set similar to that of full-parameter fine-tuning with fewer trainable parameters. In Figures 3(c) and 3(d), we compare the performance of fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA outperforms LoRA with the same amount of trainable parameters. Furthermore, as the rank increases, PiSSA quickly reaches and surpasses the performance of full-parameter fine-tuning. Notably, when the training loss of LLaMA 2-7B with a rank of 128 PiSSA equals that of full-parameter fine-tuning, it surpasses full-parameter fine-tuning in terms of generalization on downstream tasks. This validates our approach of using principal singular values and vectors to replace full-parameter model fine-tuning, ignoring components of lower rank, resulting in a model with improved robustness.
4.3 Comparative Analysis of PiSSA, LoRA, and Full Fine-Tuning Across Various Tasks
Table 1 presents a comparative evaluation of fine-tuning strategies using PiSSA, LoRA, and full parameter fine-tuning. In these experiments, we fine-tune LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [51] and MATH [2] validation sets. Additionally, the models are fine-tuned on the CodeFeedback dataset [53] and their coding proficiency is evaluated on the HumanEval [54] and MBPP [55] datasets. Furthermore, the models are trained on the WizardLM-Evol-Instruct dataset [7] and tested for their conversational abilities on the MT-Bench dataset [6].
Table 1: Performance of different models on various tasks
Model Strategy Trainable GSM8K MATH HumanEval MBPP MT-Bench
Parameters
LLaMA 2-7B Full Fine-tuning 6738M 49.05 7.22 21.34 35.59 4.91
LoRA 320M 42.3 5.5 18.29 35.34 4.58
PiSSA 320M 53.07 7.44 21.95 37.09 4.87
Mistral-7B Full Fine-tuning 7242M 67.02 18.6 45.12 51.38 4.95
LoRA 168M 67.7 19.68 43.9 58.39 4.9
PiSSA 168M 72.86 21.54 46.95 62.66 5.34
Gemma-7B Full Fine-tuning 8538M 71.34 22.74 46.95 55.64 5.4
LoRA 200M 74.9 31.28 53.66 65.41 4.98
PiSSA 200M 77.94 31.94 54.27 66.17 5.64
As can be seen from Table 1, across all models and tasks, fine-tuning with PiSSA surpasses the effects of fine-tuning with LoRA, as well as most of full fine-tuning. In the future, we will extend the comparing to more models and tasks.
Table 2: Comparation between SVD and Fast SVD in terms of initialization time, error and training loss.
Metric Niter Rank
1 2 4 8 16 32 64 128
Initialization Time 1 5.05 8.75 5.07 8.42 5.55 8.47 6.80 11.89
2 4.38 4.71 4.79 4.84 5.06 5.79 7.70 16.75
4 5.16 4.73 5.09 5.16 5.60 7.01 7.90 11.41
8 4.72 5.11 5.14 5.40 5.94 7.80 10.09 14.81
16 6.24 6.57 6.80 7.04 7.66 9.99 14.59 22.67
$\infty$ 434.92 434.15 434.30 435.42 435.25 437.22 434.48 435.84
Initialization Error 1 1.30E-3 1.33E-3 1.55E-3 1.9E-3 1.98E-3 1.97E-3 2.00E-3 1.93E-3
2 5.84E-4 1.25E-3 1.45E-3 1.43E-3 1.48E-3 1.55E-3 1.48E-3 1.33E-3
4 6.01E-4 8.75E-4 6.75E-4 1.10E-3 1.05E-3 1.03E-3 1.08E-3 9.75E-4
8 1.26E-4 2.34E-4 5.25E-4 7.25E-4 5.75E-4 8.25E-4 8.25E-4 7.75E-4
16 7.93E-5 2.25E-4 1.28E-4 6.50E-4 4.25E-4 6.50E-4 6.00E-4 4.75E-4
$\infty$
Training Loss 1 0.3629 0.3420 0.3237 0.3044 0.2855 0.2657 0.2468 0.2301
2 0.3467 0.3337 0.3172 0.2984 0.2795 0.2610 0.2435 0.2282
4 0.3445 0.3294 0.3134 0.2958 0.2761 0.2581 0.2414 0.2271
8 0.3425 0.3279 0.3122 0.2950 0.2753 0.2571 0.2406 0.2267
16 0.3413 0.3275 0.3116 0.2946 0.2762 0.2565 0.2405 0.2266
$\infty$ 0.3412 0.3269 0.3116 0.2945 0.2762 0.2564 0.2403 0.2264
4.4 Fast Singular Value Decomposition
In order to speed up the decomposition of the pre-trained matrix $W$, we adopted the algorithm proposed by Halko et.al [56] (denoted as Fast SVD), which introduces randomness to achieve an approximate matrix decomposition. We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table 4.3. Initialization time refers to the computation time taken to decompose the pre-trained parameter matrix $W$, measured in seconds. Initialization error indicates the magnitude of the discrepancy introduced by Fast SVD compared to SVD after decomposing the matrix. Specifically, the error is the sum of the absolute differences between the matrices decomposed by original SVD and Fast SVD. For the error, we report the results of the self-attention module in the table. Loss refers to the loss value at the end of training. In Fast SVD, the parameter niter refers to the number of subspace iterations to conduct. A larger niter leads to increased decomposition time but results in smaller decomposition error. The symbol $\infty$ represents the experimental results with the SVD method.
It can be observed that the computation time of the SVD is tens of times that of Fast SVD. In addition, SVD exhibits consistently high time consumption with minimal variation as the rank increases, while Fast SVD, although experiencing a slight increase in computation time with higher ranks, remains significantly lower compared to SVD throughout. As the rank increases, the initialization error initially rises gradually, with a slight decrease observed when the rank reaches 128. And at the same rank, increasing the niter in Fast SVD leads to a gradual reduction in error. For training loss, we observed that as the rank increases, the training loss decreases gradually. At the same rank, with the increase of niter, the training loss of models initialized based on Fast SVD approaches that of models initialized based on SVD.
5 Conclusion
This paper introduces a parameter-efficient fine-tuning (PEFT) technique in which singular value decomposition is applied to the weight matrix of pre-trained models to extract principal components. These components are then used to initialize an adapter named PiSSA. Fine-tuning PiSSA can closely replicate the effects of fine-tuning the complete model in the beginning, while maintaining a good parameter efficiency. Through extensive experiments, we show that PiSSA’s fine-tuning performance significantly surpasses that of LoRA. PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza. As PiSSA shares the same architecture as LoRA, it is easy to use and provides an efficient alternative initialization method. There remain several questions to be verified in the future: 1) Can PiSSA achieve significant improvements on a broader range of tasks and larger models? 2) When the iteration steps of LoRA are sufficiently long (adequately fitting the data), can it match the performance of PiSSA? 3) Can combining the successors of LoRA with PiSSA lead to further enhancement? 4) How to give a theoretical explanation for the advantages of PiSSA over LoRA? We are actively exploring these questions. Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments, and look forward to more tests and suggestions from the community.